OpenStudy (anonymous):

What is the common ratio of this geometric sequence 1 27 2 18 3 12 4 8 5 16/3 Answers: A: -9 B: 2/3 C: 3/2 D: 9

OpenStudy (anonymous):

What's the answer someone please help

OpenStudy (anonymous):

I will give you a metal someone help

OpenStudy (anonymous):

*medal

OpenStudy (anonymous):

Helppp

OpenStudy (puzzler7):

18/27 = 2/3 12/18 = 2/3 8/12 = 2/3 16/3/8 = 2/3 The answer is B: 2/3